# Exploiting Code Symmetries for Learning Program Semantics

### v7

@@ 8 Conclusion
We studied code symmetries’ impact on code LLM architectures for program reasoning tasks, introducing a novel self-attention variant that brought significant gains in generalization and robustness across a variety of program analysis tasks, providing valuable insights for specialized LLM development in reasoning and analyzing programs.
9 Broader Impact
@@ Appendix A Detailed Preliminaries
An action $\bullet$ of a symmetry group $(G,\circ)$ is a binary operation defined on a set of objects $X$, i.e., $\bullet:G\times X\rightarrow X$, [^1] In group theory literature, this is often called the left action, but we will omit “left” as it is the only type of action we will use in this paper. where
  • Identitiy: $\forall x\in X,\mathbf{1}\bullet x=x$
@@ Appendix B Complete Proofs
Theorem 2. The biased self-attention layer, $GA(e)=w_{v}e\cdot(s(w_{k}e^{T}\cdot w_{q}e)+d_{\mathcal{IG}})$, is $Aut(\mathcal{IG})$-equivariant.
  • Proof.
    $GA(\sigma\cdot e)$
    $=w_{v}\sigma(e)\cdot(s(w_{k}\sigma(e)^{T}\cdot w_{q}\sigma(e))+\sigma(d_{\mathcal{IG}}))$
    $=w_{v}ep_{\sigma}\cdot(s((w_{k}ep_{\sigma})^{T}\cdot w_{q}ep_{\sigma})+d_{\mathcal{IG}})$
    $=w_{v}ep_{\sigma}\cdot s(p_{\sigma}^{T}(w_{k}e)^{T}\cdot w_{q}ep_{\sigma})+w_{v}ep_{\sigma}\cdot d_{\mathcal{IG}}$
    $=w_{v}e(p_{\sigma}p_{\sigma}^{T})\cdot s((w_{k}e)^{T}\cdot w_{q}e)\cdot p_{\sigma}+w_{v}e\cdot d_{\mathcal{IG}}\cdot p_{\sigma}$
    $=w_{v}e\cdot((s(w_{k}e)^{T}\cdot w_{q}e)+d_{\mathcal{IG}})\cdot p_{\sigma}$
    $=\sigma(GA(e))$
@@
$=w_{v}\sigma(e)\cdot(s(w_{k}\sigma(e)^{T}\cdot w_{q}e_{i})+\sigma(d_{i}))$
@@
$=w_{v}ep_{\sigma}\cdot(s((w_{k}ep_{\sigma})^{T}\cdot w_{q}e_{i})+p^{T}_{\sigma}d_{i})$
$=w_{v}ep_{\sigma}\cdot s(p_{\sigma}^{T}(w_{k}e)^{T}\cdot w_{q}e_{i})+w_{v}ep_{\sigma}\cdot p^{T}_{\sigma}d_{i}$
$=w_{v}e(p_{\sigma}p_{\sigma}^{T})\cdot s((w_{k}e)^{T}\cdot w_{q}e_{i})+w_{v}e\cdot(p_{\sigma}p^{T}_{\sigma})\cdot d_{i}$
@@ D.1 Implementation Details
We implement SymC using Fairseq (Ott et al., 2019) PyTorch (Paszke et al., 2019). To compute PDG for x86 assembly code, we utilize Ghidra to lift the assembly code into P-Code, an intermediate representation used by Ghidra, to track implicit data and control flow via the FLAGS register. To compute PDG for Java functions, we employ JavaParser on Java ASTs to analyze control and data dependencies. We conduct all the experiments on three Linux servers with Ubuntu 20.04 LTS, each featuring an AMD EPYC 7502 processor, 128 virtual cores, and 256GB RAM, with 12 Nvidia RTX 3090 GPUs in total.
Datasets. We use the Java dataset collected by Allamanis et al. (2016) to evaluate the function name prediction. The dataset includes 11 Java projects, such as Hadoop, Gradle, etc., totaling 707K methods and 5.6M statements. We fix Hadoop as our test set and use the other projects for training, to ensure the two sets do not overlap. For binary analysis, we collect and compile 27 open-source projects such as OpenSSL, which contains approximately 1.13M functions and 137.6M instructions.
@@ D.2 Experiment Configurations
Evaluation metrics. For most analysis tasks (§5), we use F1 score, the harmonic mean of precision and recall. For function similarity detection, we use Area Under Curve (AUC) of ROC curve, as it handles continuous similarity scores and varying thresholds for determining function similarity. We note that AUC-ROC might not be the most reliable metric (Arp et al., 2022), but we choose it primarily for comparing to the baselines whose results are measured in AUC-ROC (Li et al., 2021).
@@ Appendix E Additional Experiments
Function Name SymC 68.4M 202M 0.363 $*{}^{*}$ 0.363 0.363 0.363 0 $*{}^{*}$ 0 0
@@
$*{}^{*}$e observe a slight value change due to the floating point precision error by adopting memory-efficient 16-bit.
@@ E.2 Efficiency
Cost of graph computation. To incorporate the inductive bias of code, many code models involve extracting and encoding code structures, including our baselines, e.g., GraphCodeBERT and GGNN. This is because computing PDGs statically is not overly expensive. Table 7 shows the runtime performance (in milliseconds) per code sample of SymC and PalmTree on extracting code structures, training, and inference, using the same exact hardware, i.e., an AMD EPYC 7502 processor, 128 virtual cores, 256GB RAM, Nvidia RTX 3090 GPUs.
Table 7: The speed (in seconds) of building code structures, i.e., in our case is building PDG, training, and inference for SymC and PalmTree per code sample.
Graph Construction Training Inference
@@
PalmTree 89.67 31.38 11.64
SymC’s cheap computation of PDG incurs 17.8$\times$ less runtime overhead than PalmTree. However, our approach does incur additional computational cost for graph construction. Therefore, it remains an interesting research problem to incorporate system optimization, e.g., caching, to improve the efficiency of the PDG computation during inference.
[…]
SymC 0.07 0.025 0.009
$*{}^{*}$ 89.67 31.38 11.64
$*{}^{*}$almTree did not disclose its hours for pre-training, so we include the

### v9

@@ 9 Conclusion
We studied code symmetries’ impact on code LLM architectures for program reasoning. We introduced a novel self-attention variant that guarantees equivariance against code symmetries based on the permutation group. We implement SymC and evaluate against various semantics-preserving transformations across different program analysis tasks, demonstrating the improved generalization in reasoning and analyzing programs.
Acknowledgement
We thank the anonymous reviewers for their constructive comments and feedback, which significantly improved this paper. This work was supported in part by NSF grants CNS-2154874, CNS-1564055, ONR grant N00014-17-1-2788, an NSF career award, a Google ASPIRE Award, multiple Google Cyber NYC awards, Columbia SEAS/EVPR Stimulus award, Columbia SEAS-KFAI Generative AI and Public Discourse Research award, and Accenture.
Impact Statement
@@ Appendix A Detailed Preliminaries
An action $\bullet$ of a symmetry group $(G,\circ)$ is a binary operation defined on a set of objects $X$, i.e., $\bullet:G\times X\rightarrow X$ [^1] In group theory literature, this is often called the left action, but we will omit “left” as it is the only type of action we will use in this paper., where
  • Identity: $\forall x\in X,\mathbf{1}\bullet x=x$
@@ Appendix B Complete Proofs
$=w_{v}\sigma(e)\cdot s(w_{k}\sigma(e)^{T}\cdot w_{q}e_{i}+\sigma(d_{i}))$
@@
$=w_{v}ep_{\sigma}\cdot s((w_{k}ep_{\sigma})^{T}\cdot w_{q}e_{i}+p^{T}_{\sigma}d_{i})$
$=w_{v}ep_{\sigma}\cdot s(p_{\sigma}^{T}(w_{k}e)^{T}\cdot w_{q}e_{i}+p^{T}_{\sigma}d_{i})$
$=w_{v}e(p_{\sigma}p_{\sigma}^{T})\cdot s((w_{k}e)^{T}\cdot w_{q}e_{i}+d_{i})$
@@
Based on Lemma 2 and Lemma 3, we now prove Theorem 2 – the biased self-attention layer is $Aut(\mathcal{IG})$-equivariant.
  • Proof.
    $GA(\sigma\cdot e)$
    $=w_{v}\sigma(e)\cdot s(w_{k}\sigma(e)^{T}\cdot w_{q}\sigma(e)+\sigma(d_{\mathcal{IG}}))$
    $=w_{v}ep_{\sigma}\cdot s((w_{k}ep_{\sigma})^{T}\cdot w_{q}ep_{\sigma}+d_{\mathcal{IG}})$
    $=w_{v}e(p_{\sigma}p_{\sigma}^{T})\cdot s((w_{k}e)^{T}\cdot w_{q}e\cdot p_{\sigma}+d_{\mathcal{IG}}\cdot p_{\sigma})$
    $=w_{v}e\cdot s((w_{k}e)^{T}\cdot w_{q}e+d_{\mathcal{IG}})\cdot p_{\sigma}$
    $=\sigma(GA(e))$
@@ D.1 Implementation Details
We implement SymC using Fairseq (Ott et al., 2019) PyTorch (Paszke et al., 2019). We conduct all the experiments on three Linux servers with Ubuntu 20.04 LTS, each featuring an AMD EPYC 7502 processor, 128 virtual cores, and 256GB RAM, with 12 Nvidia RTX 3090 GPUs in total.
PDG construction. To compute PDG for x86 assembly code, we utilize Ghidra to lift the assembly code into P-Code, an intermediate representation used by Ghidra, to track implicit data and control flow via the FLAGS register. The key advantage of using Ghidra P-Code is that it keeps the side effects of the instructions, e.g., manipulating flag registers, explicit. For example, cmp instruction will set the zero flag implicitly at the assembly level, but Ghidra P-Code will translate it into a series of IR instructions with one of them operating on the ZF explicitly. We then analyze the data and control dependencies between each pair of P-Code instructions and flag the corresponding assembly code pair as dependent if at least one dependent P-Code instruction pair exists between those of the assembly code pair.
To compute PDG for Java functions, we employ JavaParser on Java ASTs for each statement to analyze control and data dependencies. We iterate through every pair of the statements and connect the dependent statement pairs with the directed edge. We extract control-flow dependencies between statements by connecting edges between different basic blocks to prevent permutations among basic blocks.
Datasets. We use the Java dataset collected by Allamanis et al. (2016) to evaluate the function name prediction. The dataset includes 11 Java projects, such as Hadoop, Gradle, etc., totaling 707K methods and 5.6M statements. We fix Hadoop as our test set and use the other projects for training, to ensure the two sets do not overlap.
For binary analysis, we collect and compile 27 open-source projects, such as OpenSSL, ImageMagic, CoreUtils, SQLite, etc., which contain approximately 1.13M functions and 137.6M instructions. We categorize the binaries based on the compilers (GCC or Clang), optimizations (O0-O3), and obfuscations (using a LLVM-based obfuscation passes based on Hikari (Zhang, 2017)) and show their statistics in Table 4.
Table 4: The statistics of our binary dataset, categorized by compilers, optimizations, obfuscations, and lengths.
# Files # Functions # Instructions
Different Compilers
GCC 1,140 274,840 33,464,420
Clang 1,136 279,832 31,949,673
Different Optimization Levels
O0 285 98,451 10,202,328
O1 285 61,298 7,096,903
O2 285 61,298 7,096,903
O3 285 57,023 9,101,578
Different Compiler Obfuscations
bcf 158 61,701 9,173,168
cff 158 59,724 11,146,990
ind 158 56,291 2,501,422
spl 158 61,379 9,652,268
sub 158 59,694 6,198,900
@@ D.2 Experiment Configurations
Program analysis tasks for evaluation. For source code analysis tasks, we focus on the method name prediction and defect prediction. Method name prediction aims to predict the function name (in natural language) given the body of the method. This task has been extensively evaluated by prior works to test the generalizability of code models (Rabin et al., 2021). Following the strategies adopted in SymLM (Jin et al., 2022), we tokenize the function names and formulate the function name prediction as a multi-label classification problem, i.e., multiple binary classifications that predict the presence of a specific token in the vocabulary. We then match the predicted tokens with the tokenized ground truth tokens to compute the F1 score. We thus employ a 2-layer fully-connected network $F:\mathbb{R}^{d}\rightarrow\{0,1\}^{L}$ on top of a mean-pooled embedding from self-attention layers to ensure $Aut(PDG)$-invariance (§3.5), where $L$ is the vocabulary of all function name tokens in our dataset.
The defect prediction task is much more simplified than method name prediction. It is a binary classification task to predict whether a given Java method is buggy or not. We obtain the dataset from Defects4J (Just et al., 2014).
We consider three binary analysis tasks commonly used to evaluate ML-based approaches to security applications (Li et al., 2021). The first task is function similarity detection. It aims to detect semantically similar functions, e.g., those compiled by different compiler transformations (see below). This task is often used to detect vulnerabilities, i.e., by searching similar vulnerable functions in firmware, or malware analysis, i.e., by searching similar malicious functions to identify the malware family (Marcelli et al., 2022; Xu et al., 2017). We leverage the pooling-based predictor (§3.5) by taking the mean of the embeddings $e$ produced by the last self-attention layer and feed that to a 2-layer fully-connected neural network $F:\mathbb{R}^{d}\rightarrow\mathbb{R}^{d}$. We then leverage the cosine distance between the output of $F$ for a pair of function embeddings, i.e., $e^{1},e^{2}$, to compute their similarity: $cos(F(\mu(e^{1})),F(\mu(e^{2})))$.
The second task is function signature prediction (Chua et al., 2017). It aims to predict the number of arguments and their source-level types given the function in stripped binaries. Similar to function similarity detection, we stack a 2-layer fully-connected network $F:\mathbb{R}^{d}\rightarrow L$ on top of mean-pooled embeddings from self-attention layers, which outputs the function signature label, e.g., $L=\{\texttt{int},\texttt{float},...\}$.
[…]
@@
Evaluation metrics. For most analysis tasks (§5), we use F1 score, the harmonic mean of precision and recall. We follow the existing works (Jin et al., 2022) and adopt their definition of F1 beyond the binary classifier. Take function name prediction as an example, we first tokenize both the ground truth and the predicted function names into a set of tokens, i.e., $W$ and $W^{\prime}$, respectively. In this case, precision measures out of $W^{\prime}$, how many tokens in $W^{\prime}$ appear in $W$: $precision=\frac{W\cap W^{\prime}}{|W^{\prime}|}$, and recall measures out of all the tokens in $W$, how many of them are correctly predicted in $W^{\prime}$: $recall=\frac{W\cap W^{\prime}}{|W|}$. We measure the precision and recall and compute the F1 score for each sample accordingly. We then average them across all samples.
For function similarity detection, as the cosine distance between two function embeddings can be an arbitrary real value between -1 and 1, a threshold is needed to determine whether pairs are similar or not. Therefore, we employ the ROC curve by varying the thresholds and measuring the corresponding True Positive Rate (TPR): TPR=True PositivesTrue Positives+False $\text{TPR}=\frac{\text{True Positives}}{\text{True Positives}+\text{False Negatives}}$ and False Positive Rate (FPR): FPR=False PositivesFalse Positives+True $\text{FPR}=\frac{\text{False Positives}}{\text{False Positives}+\text{True Negatives}}$. The ROC curve is then plotted with FPR at the x-axis and TPR at the y-axis. Following Li et al. (2021), we leverage the Area Under Curve (AUC) score of the ROC curve to quantify the performance for ease of comparison.
We note that AUC-ROC might not be the most reliable metric (Arp et al., 2022), but we choose it primarily for comparing to the baselines whose results are measured in AUC-ROC (Li et al., 2021).
@@ Appendix E Additional Experiments
Function Name SymC 68.4M 202M 0.363 0.364 0.363 0.363 0.363 0 0.1 0 0
@@
We observe a slight value change due to the floating point precision error by adopting memory-efficient 16-bit.
@@ E.2 Efficiency
Overhead. To incorporate the inductive bias of code, many code models involve extracting and encoding code structures, including our baselines, e.g., GraphCodeBERT and GGNN. This is because computing PDGs statically is not overly expensive. Figure 11 shows the runtime performance (in milliseconds) per code sample of SymC and PalmTree on extracting code structures, training, and inference, using the same exact hardware.
Figure 11: Comparing SymC to standard Transformer encoder in terms of the additional overhead introduced by constructing PDG, and training and inference with PDG-biased self-attention layers.
SymC’s cheap computation of PDG incurs 88.8$\times$ less runtime overhead than PalmTree. However, our approach does incur additional computational cost for graph construction. Therefore, it remains an interesting research problem to incorporate system optimization, e.g., caching, to improve the efficiency of the PDG computation during inference.
Training efficiency. We study the training effort (including both pre-training and fine-tuning) of SymC and PalmTree. Table 8 shows their GPU hours, power, and emitted carbon dioxide estimation when they reach 0.5 F1 score in memory region prediction. We assume the GPU always reaches its power cap (350W) to estimate an upper bound of the power usage. CO2eq stands for the carbon dioxide equivalent, a unit for measuring carbon footprints. By being more training efficient, SymC incurs 1,281$\times$ less total GPU time, power, and emitted carbon dioxide than PalmTree in obtaining the same performance.
Time (Hours) Power (kWh) Carbon (CO2eq)
@@
PalmTree-O 89.67 31.38 11.64
PalmTree did not disclose its hours for pre-training, so we include the